Radical 2y-5 = 1 + Radical y-3
Answer by nyc_function(2741)

sqrt{2y - 5} = 1 + sqrt{y - 3}
Isolate the radicals.
sqrt{2y - 5} - sqrt{y - 3} = 1
Square both sides.
[sqrt{2y - 5} - sqrt{y - 3}]^2 = (1)^2
Can you take it from here?